OpenStudy (anonymous):

Translate the following situation into an equation. DO NOT SOLVE.. Suppose a crate of oranges weighs eight pound more than a crate of banana's and a crate of kiwi weighs one-third as much as a crate of banana's. Together the 3 crates weigh 43 pounds. What is the weight of the crate of banana's.

OpenStudy (anonymous):

O = B+8 3K=B O+B+K=43

OpenStudy (anonymous):

O for oranges B for bananas K for kiwis

OpenStudy (anonymous):

sorry i have little imagination

OpenStudy (anonymous):

That is ok

OpenStudy (anonymous):

clear though? can you account for each equation?

OpenStudy (anonymous):

so the final equation you are looking for is B+B+8+B/3 = 43